# Dynamic Decoupling in Multidimensional Screening

### v1

@@ 3.1 Additional Periods
If the optimal mechanism is to simply sell goods independently, do dynamics have bite? The answer is no. Consider an alternative environment where instead of buying $n$ products simultaneously, there were $n+1$ periods. In the first period, the agent draws $\gamma$ and the monopolists writes a contract. In each subsequent period $i$ the buyer’s valuation $\theta^{i}$ is drawn according to $F^{i}(\theta^{i}|\gamma)$ and good $i$ is sold via some mechanism. Call this the sequential case. In general, profits from the original problem of selling simultaneously are upper bounded by profits from the alternative case of selling sequentially: When selling simultaneously, the buyer draws all of their valuations at once and is able to better coordinate joint deviations across goods. As such, any incentive compatible mechanism in the original problem is also incentive compatible in the sequential case, implying that the value of the sequential case is lower-bounded by the value of the original problem. However, this bound is tight whenever values are independent: Endowing the designer with additional power to screen intertemporally does not increase profits. Since all information rents are pulled forward into period one anyways, additional periods beyond period two are inconsequential. Note that this result is not a direct consequence of EsoSzentes2017: Dynamic irrelevance in their setting only pertains to not needing to pay information rents to elicit future information so the same allocation rule leads to the same profit. I contribute by showing that even when expanding the space of implementable allocation rules, the principle does not do better when screening additionally through time. As such, the role of $\gamma$ (learning about learning) is qualitatively different than $\theta$ (learning about valuations).
@@
If there is dependence and different dimensions are drawn sequentially, however, different components of $\theta$ leads to both learning about learning and learning about valuations. The proof of Proposition 1 compares the solution to a multidimensional sequential screening problem to the optimal mechanism identified in segal2014; absent independence, the optimal mechanism in the sequential case is difficult to derive.
@@ 3.2 Observed Orthogonalization
The standard approach to solving dynamic contracting problems is via orthogonalization, transforming the agent’s future information so that it is independent of their current information. In this case, EsoSzentes2017 shows dynamic irrelevance in a different sense: For any implementable allocation rule, making the agent’s future orthogonalization information publicly observable does not change the monopolist’s profits. Does that result hold in the multidimensional setting and can it be operationalized?
@@
By similar reasoning as Lemma 1, we are able to recover the irrelevance result of EsoSzentes2017, even when there is dependence between dimensions (recall that Lemma 1 did not require independence either). However, Lemma 1 clarifies the mechanism by which post-contractual information rents need not be paid: they have simply already been extracted in period one. On the other hand, the reasoning of EsoSzentes2017 directly jumps to the case where no information rents need to be paid.
@@
If $F$ has invariant dependencies, then the solution of the original problem coincides with the solution of the relaxed problem.
However, invariant dependencies is violated, solutions to the relaxed problem are not implementable in the original problem and cyclic monotonicity becomes the key binding constraint; in a sequential sequential problem, cyclic monotonicity does not have any bite as contracting is done one dimension at a time.
@@ 4 Discussion
Endowing a principle the ability to contract intertemporally enables frontloading surplus extraction. This distorts the information rents which need to be paid to the agent in favor of the principle, increasing profits and simplifying the multidimensional screening problem when distributions satisfy invariant dependencies. In this case, the distortion of information rents enables the monopolist’s problem to be converted into one that is separable across dimensions; this transformation solves a probability continuity equation and corresponding to statistical tests of the agent’s reported pre-contractual information being separable across dimensions. When this is the case, the optimal mechanism sells goods separately, in contrast to complicated non-linear bundling usually featured in multidimensional screening problems. However, this mechanism parallels membership schemes retailers often employ in practice. Finally, these results formalize new notions of dynamic irrelevance in dynamic contracting.
@@
There are several avenues for further investigation. First, offering the optimal sequential screening mechanism independently for each good is still a feasible mechanism even when invariant dependencies does not hold: Are there more general conditions for which it is optimal? Away from the invariant dependencies case, the function $\mathbf{V}$ may no longer be the same. However, a more complicated $\mathbf{V}$ may still be independent across dimensions, leading to a different pointwise maximization solution; even if the pointwise maximum is not implementable due to violations of cyclic monotonicity, it may still be the case that contracting independently across goods remains optimal. Second is analyzing the welfare effects of endowing the principal the ability to screen through time. Compared to the case where the monopolist can only contract after valuations themselves are known to the consumer, the monopolist always makes more profits in the sequential case (ignoring the agent’s pre-contractual private information and offering the optimal static multidimensional screening mechanism is always feasible). On the other hand, consumer information rents are distorted downwards but since it costs the monopolist less to procure information rents, realized information rents (net of distortions) may be larger. Investigating both how total surplus changes and how the division of surplus changes would be of interest.
@@
Thus, $t_{1},t_{2},q$ are all pinned down by $u$ so the designer only needs to maximize over $u$; the monopolist’s objective function can be written as
@@
$=\sum_{j}\left[f\left(\frac{f^{j}f_{\gamma}^{j}-F_{\gamma}^{j}f^{j}_{\theta_{j}}}{(f^{j})^{2}}\right)+\frac{F_{\gamma}^{j}}{f^{j}}f\frac{\partial}{\partial\theta^{j}}\left(\ln(c(F^{1},...,F^{n}))+\sum_{\ell}\ln(f^{\ell})\right)\right]$
$=f\sum_{j}\left[\left(\frac{f^{j}f_{\gamma}^{j}-F_{\gamma}^{j}f^{j}_{\theta_{j}}}{(f^{j})^{2}}\right)+\frac{F_{\gamma}^{j}}{f^{j}}\frac{\partial\ln(c)}{\partial u^{j}}f^{j}+F_{\gamma}^{j}\frac{f^{j}_{\theta^{j}}}{(f^{j})^{2}}\right]$
@@
Thus, $q(\gamma,\theta)$ is cyclically monotone in $\theta$ for each $\gamma$ and thus induces a utility function $u(\gamma,\theta)$ that is 1-Lipschitz and convex in $\theta$. As virtual values are increasing in $\gamma$, the threshold $\hat{\theta^{j}}(\gamma)$ is decreasing in $\gamma$ so $q^{j}(\gamma,\theta)$ is increasing in $\gamma$. Thus, first-period incentive constraints hold as well. ∎
@@
Let $t(\gamma,\theta)$ be net transfers. First, any $q(\gamma,\theta),t(\gamma,\theta)$ that is implementable in the original setting is implementable in the sequential setting. We will prove the contrapositive: If $q(\gamma,\theta),t(\gamma,\theta)$ is not implementable in the sequential setting, then it must not be implementable in the original setting.
Suppose $q(\gamma,\theta),t(\gamma,\theta)$ is not implementable in the sequential setting. Then, the buyer must have a profitable deviation at some point along the path of play of the mechanism. For notational simplicity, suppose the deviation happens after $\gamma$ has been truthfully reported; the same reasoning holds for deviations in $\gamma$ as well. Thus, there exists some $\gamma,\theta^{1},...,\theta^{k}$ such that the buyer has learned $\gamma,\theta^{1},...,\theta^{k}$, truthfully reported $\gamma,\theta^{1},...,\theta^{k-1}$, and can profitably deviate to a sequence of misreports $\{\tilde{\theta}^{j}(\theta_{1},...,\theta_{j})\}_{j=k}^{n}$. Slightly abusing notation, let $\tilde{\theta}(\theta)$ be the agent’s reports for a given realization of $\theta$. For this to be profitable, we have
@@
Then, the monopolist’s objective in the sequential case, as formalized in segal2014, coincides with the monopolist’s objective in the original setting, albeit with different constraints. Since pointwise maximization is implementable in the original problem, it is also implementable in the sequential case. ∎
@@
$\mathbb{E}_{\gamma}\left[\hat{t}(\gamma)\right]=\mathbb{E}_{\gamma,z}\left[\hat{q}(\gamma,z)\cdot\left(v(\gamma,z)+v_{\gamma}(\gamma,z)\frac{1-G(\gamma)}{g(\gamma)}\right)\right].$
@@
$\mathbb{E}_{\gamma}\left[\hat{t}(\gamma)\right]=\mathbb{E}_{\gamma,\theta}\left[q(\gamma,\theta)\cdot\left(\theta+\mathbf{V}(\theta|\gamma)\frac{1-G(\gamma)}{g(\gamma)}\right)\right].$

### v2

@@ 5.1 Additional Periods
If the optimal mechanism is to simply sell goods independently, do dynamics have bite? The answer is no. Consider an alternative environment where instead of buying $n$ products simultaneously, there were $n+1$ periods. In the first period, the buyer draws $\gamma$ and the monopolists writes a contract. In each subsequent period $i$ the buyer’s valuation $\theta^{i}$ is drawn according to $F^{i}(\theta^{i}|\gamma)$ and good $i$ is sold via some mechanism. Call this the sequential case. In general, profits from the original problem of selling simultaneously are upper bounded by profits from the alternative case of selling sequentially: When selling simultaneously, the buyer draws all of their valuations at once and is able to better coordinate joint deviations across goods. As such, any incentive compatible mechanism in the original problem is also incentive compatible in the sequential case, implying that the value of the sequential case is lower-bounded by the value of the original problem. However, this bound is tight whenever values are independent: Endowing the monopolist with additional power to screen intertemporally does not increase profits. Since all information rents are pulled forward into period one anyways, additional periods beyond period two are inconsequential. Note that this result is not a direct consequence of Eső and Szentes (2017): Dynamic irrelevance in their setting only pertains to not needing to pay information rents to elicit future information so the same allocation rule leads to the same profit. I contribute by showing that even when expanding the space of implementable allocation rules, the principle does not do better when screening additionally through time. As such, the role of $\gamma$ (learning about learning) is qualitatively different than $\theta$ (learning about valuations).
@@
If there is dependence and different dimensions are drawn sequentially, however, different components of $\theta$ leads to both learning about learning and learning about valuations. The proof of Proposition 2 compares the solution to a multidimensional sequential screening problem to the optimal mechanism identified in Pavan et al. (2014); absent independence, the optimal mechanism in the sequential case is difficult to derive.
@@ 5.2 Observed Orthogonalization
The standard approach to solving dynamic contracting problems is via orthogonalization, transforming the buyer’s future information so that it is independent of their current information. In this case, Eső and Szentes (2017) shows dynamic irrelevance in a different sense: For any implementable allocation rule, making the buyer’s future orthogonalization information publicly observable does not change the monopolist’s profits. Does that result hold in the multidimensional setting and can it be operationalized?
@@
By similar reasoning as Lemma 1, we are able to recover the irrelevance result of Eső and Szentes (2017), even when there is dependence between dimensions (recall that Lemma 1 did not require independence either). However, Lemma 1 clarifies the mechanism by which post-contractual information rents need not be paid: they have simply already been extracted in period one. On the other hand, the reasoning of Eső and Szentes (2017) directly jumps to the case where no information rents need to be paid.
@@
If assumptions A1-A3 are satisfied, then the solution of the original problem coincides with the solution of the relaxed problem.
However, if invariant dependencies is violated, solutions to the relaxed problem are not implementable in the original problem and cyclic monotonicity becomes the key binding constraint. In the sequential version of the problem, cyclic monotonicity does not have any bite as contracting is done one dimension at a time.
5.3 Downstream Auction
Suppose there were $B$ bidders indexed by subscripts with generic element $b$. Each bidder observes $\gamma_{b}\sim G_{b}(\gamma)$ before contracting with the auctioneer and later observes $\theta_{b}\sim F_{b}(\theta_{b}|\gamma_{b})$. [^6] As in Eső and Szentes (2007), we could also allow the auctioneer to observe post-contractual information via the multidimensional orthogonalization developed in Section 5.2. The auctioneer chooses an allocation rule
$q:\prod_{b}\Gamma_{b}\times\prod_{b}\Theta_{b}\to[0,1]^{N\times B}$
mapping all private information to quantities of each good for each bidder. Following the same processes as this paper and Eső and Szentes (2007), the auctioneer’s problem can be written as:
$\max_{q(\{\gamma_{b},\theta_{b}\}_{b})}\text{ }$ $\mathbb{E}_{\{\gamma_{b},\theta_{b}\}_{b}}\left[\sum_{j}q^{j}_{b}(\gamma_{b},\theta_{b})\phi^{j}_{b}(\gamma_{b},\theta^{j}_{b})\right]$
[…]
$q_{b}(\gamma_{b},\theta_{b})\in[0,1]^{n}\text{ for all }\gamma,\theta,b;$
$\sum_{b}q^{j}_{b}(\{\gamma_{b},\theta_{b}\}_{b})\leq 1\text{ for each }j.$
$IC1,IR.$
where
$\phi^{j}_{b}(\gamma_{b},\theta^{j}_{b})=\left(\theta^{j}_{b}+\frac{F_{\gamma,b}^{j}(\theta^{j}_{b}|\gamma_{b})}{f^{j}_{b}(\theta^{j}_{b}|\gamma_{b})}\frac{1-G_{b}(\gamma_{b})}{g_{b}(\gamma_{b})}\right)$
denotes bidder $b$’s virtual value for good $j$.
The pointwise solution to the auctioneer’s problem is to set
$q^{j}_{\beta}(\{\gamma_{b},\theta_{b}\}_{b})=\begin{cases}1&\text{if}\;\phi^{j}_{\beta}(\gamma_{\beta},\theta^{j}_{\beta})\geq 0,\phi^{j}_{\beta}(\gamma_{\beta},\theta^{j}_{\beta})\geq\phi^{j}_{b}(\gamma_{b},\theta^{j}_{b})\text{ for all }b\\ 0&\text{else}.\end{cases}$
That is, for each good sell to the bidder with the highest non-negative virtual value and do not sell otherwise. Since virtual values are monotone in $\gamma_{b}$ for each bidder $b$ and good $j$ (by A2), this allocation rule is monotone in $\gamma_{b},\theta^{j}_{b}$ and satisfies global period one incentive constraints (by A1). Since virtual values are monotone in $\theta^{j}_{b}$ (by A2) and the allocation rule is additive across goods, cyclic monotonicity is satisfied. Since these properties hold pointwise for any realization of other bidders’ private information, taking expectations over b-b gives that all constraints are satisfied in the aggregate. This allocation rule can be implemented by running a handicap auction for each good independently of other goods.
@@ 6 Discussion
Endowing a principle the ability to contract intertemporally enables frontloading surplus extraction. This distorts the information rents which need to be paid to the buyer in favor of the principle, increasing profits and simplifying the multidimensional screening problem when distributions satisfy invariant dependencies. In this case, the distortion of information rents enables the monopolist’s problem to be converted into one that is separable across dimensions; this transformation solves a probability continuity equation and corresponds to a statistical test of the buyer’s reported pre-contractual information being separable across dimensions. When this is the case, the optimal mechanism sells goods separately, in contrast to complicated non-linear bundling usually featured in multidimensional screening problems. However, this mechanism parallels membership schemes retailers often employ in practice. Finally, these results formalize new notions of dynamic irrelevance in dynamic contracting.
@@
There are several avenues for further investigation. First, offering the optimal sequential screening mechanism independently for each good is still a feasible mechanism even when invariant dependencies does not hold: Are there more general conditions for which it is optimal? Away from the invariant dependencies case, the function $\mathbf{V}$ may no longer be the same. However, a more complicated $\mathbf{V}$ may still be independent across dimensions, leading to a different pointwise maximization solution; even if the pointwise maximum is not implementable due to violations of cyclic monotonicity, it may still be the case that contracting independently across goods remains optimal. In a similar vein, if the monopolist had robustness concerns about the worst-case dependency structure, separate sales may once again be optimal. Second is analyzing the welfare effects of endowing the principal the ability to screen through time. Compared to the case where the monopolist can only contract after valuations themselves are known to the consumer, the monopolist always makes more profits in the sequential case (ignoring the buyer’s pre-contractual private information and offering the optimal static multidimensional screening mechanism is always feasible). On the other hand, consumer information rents are distorted downwards but since it costs the monopolist less to procure information rents, realized information rents (net of distortions) may be larger. Investigating both how total surplus changes and how the division of surplus changes would be of interest. Third, Ball et al. (2026) considers multiple (single-good) firms that engage in sequential screening competing against one another. Their analysis follows a similar approach of considering the relaxed problem and finding equilibria before verifying global incentive constraints. Perhaps their results extend straightforwardly to a multi-good setting as the relaxed problem becomes separable cross dimensions. In this world, several new questions can be considered; under the optimal mechanism in this paper, upfront transfers were good-by-good but presented as aggregated. If there were multiple firms, consumers may want to only pay the upfront fee for some subset of goods at different monopolists. However, bundling upfront payments for different goods may allow firms to lock in consumers. For example, buyers cannot just buy a membership to Costco’s produce section; as a result of buying a Costco membership they purchase other goods in addition to the produce they came for. This also discourages the consumer from purchasing a membership at Costco’s competitors. Finally, many real-world sequential screening mechanisms feature a finite set of upfront and eventual strike prices, whereas the optimal mechanisms derived in this paper generally vary continuously and thus induce menus that have an uncountably infinite number of upfront and eventual strike prices. Characterizing optimal mechanisms when the menu size is restricted may further rationalize the pricing schemes seen across many industries.
@@
Thus, $t_{1},t_{2},q$ are all pinned down by $u$ so the monopolist only needs to maximize over $u$; the monopolist’s objective function can be written as
@@
$=\sum_{j}\left[f\left(\frac{f^{j}f_{\gamma}^{j}-F_{\gamma}^{j}f^{j}_{\theta^{j}}}{(f^{j})^{2}}\right)+\frac{F_{\gamma}^{j}}{f^{j}}f\frac{\partial}{\partial\theta^{j}}\left(\ln(c(F^{1},...,F^{n}))+\sum_{\ell}\ln(f^{\ell})\right)\right]$
$=f\sum_{j}\left[\left(\frac{f^{j}f_{\gamma}^{j}-F_{\gamma}^{j}f^{j}_{\theta^{j}}}{(f^{j})^{2}}\right)+\frac{F_{\gamma}^{j}}{f^{j}}\frac{\partial\ln(c)}{\partial u^{j}}f^{j}+F_{\gamma}^{j}\frac{f^{j}_{\theta^{j}}}{(f^{j})^{2}}\right]$
@@
This is also the price $p^{j}(\gamma)$ a buyer reporting $\gamma$ faces when deciding whether or not to (eventually) purchase good $j$. Then, $q(\gamma,\theta)$ is cyclically monotone in $\theta$ for each $\gamma$ and thus induces a utility function $u(\gamma,\theta)$ that is 1-Lipschitz and convex in $\theta$. As virtual values are increasing in $\gamma$, the threshold $\hat{\theta^{j}}(\gamma)$ is decreasing in $\gamma$ so $q^{j}(\gamma,\theta)$ is increasing in $\gamma$. Using divergence theorem a la Lemma 2, we can write
$U_{\gamma}(\gamma^{\prime}|\gamma)=\int_{\Theta}u(\gamma,\theta)f_{\gamma}(\theta|\gamma)d\theta=\int_{\Theta}q(\gamma,\theta)\cdot[-F_{\gamma}(\theta|\gamma)]d\theta.$
As long as $q$ is increasing in $\gamma$, Assumption A1 gives that $F_{\gamma}(\theta|\gamma)\geq 0$ so $U(\gamma^{\prime}|\gamma)$ has increasing differences and thus global incentive constraints are satisfied. Next, integrating both sides and using separability of $q$ gives
$U(\gamma)=\sum_{j}\int_{\underline{\gamma}}^{\gamma}\int_{\Theta}q(\gamma,\theta)[-F_{\gamma}^{j}(t|\gamma)]dtd\gamma=\sum_{j}\int_{\underline{\gamma}}^{\gamma}\int_{p^{j}(\gamma)}^{\bar{\theta}^{j}}-F_{\gamma}^{j}(t|\gamma)dtd\gamma$
while total expected gains from trade realized from a buyer with type $\gamma$ is
$\int_{\Theta}u(\gamma,\theta)dF(\theta|\gamma)=\mathbb{E}_{\theta|\gamma}\left[\sum_{j}\max\{\theta^{j}-p(\gamma),0\}\right]=\sum_{j}\int_{p^{j}(\gamma)}^{\bar{\theta}^{j}}1-F^{j}(\theta^{j}|\gamma)d\theta^{j}.$
Period one transfers is equal to total surplus minus consumer surplus, so
$t_{1}(\gamma)$ $=\sum_{j}\int_{p^{j}(\gamma)}^{\bar{\theta}^{j}}1-F^{j}(\theta^{j}|\gamma)d\theta^{j}-\left[\sum_{j}\int_{\underline{\gamma}}^{\gamma}\int_{p^{j}(\gamma)}^{\bar{\theta}^{j}}-F_{\gamma}^{j}(t|\gamma)dtd\gamma\right]$
$=\sum_{j=1}^{n}\left[\int_{p^{j}(\gamma)}^{\bar{\theta}^{j}}1-F^{j}(\theta^{j}|\gamma)d\theta^{j}+\int_{\underline{\gamma}}^{\gamma}\int_{p^{j}(\gamma)}^{\bar{\theta}^{j}}F_{\gamma}^{j}(t|\gamma^{\prime})dtd\gamma^{\prime}\right].$
@@
Let $t(\gamma,\theta)$ be net transfers. First, any $q(\gamma,\theta),t(\gamma,\theta)$ that is implementable in the original setting is implementable in the sequential setting. [^7] Suppose that goods are allocated at the end but information must be reported as soon as it arrives, so the allocation of good $j$ could, in theory, depend on the buyer’s reported value of $\theta^{k}$ even if $k>j$. We will prove the contrapositive: If $q(\gamma,\theta),t(\gamma,\theta)$ is not implementable in the sequential setting, then it must not be implementable in the original setting.
Suppose $q(\gamma,\theta),t(\gamma,\theta)$ is not implementable in the sequential setting. Then, the buyer must have a profitable deviation at some point along the path of play of the mechanism. For notational simplicity, suppose the deviation happens after $\gamma$ has been truthfully reported; the same reasoning holds for deviations in $\gamma$ as well. Thus, there exists some $\gamma,\theta^{1},...,\theta^{k}$ such that the buyer has learned $\gamma,\theta^{1},...,\theta^{k}$, truthfully reported $\gamma,\theta^{1},...,\theta^{k-1}$, and can profitably deviate to a sequence of misreports $\{\tilde{\theta}^{j}(\theta^{1},...,\theta^{j})\}_{j=k}^{n}$. Slightly abusing notation, let $\tilde{\theta}(\theta)$ be the buyer’s reports for a given realization of $\theta$. For this to be profitable, we have
@@
Then, the monopolist’s objective in the sequential case, as formalized in Pavan et al. (2014), coincides with the monopolist’s objective in the original setting, albeit with different constraints. Since pointwise maximization is implementable in the original problem, it is also implementable in the sequential case. ∎
@@
$\mathbb{E}_{\gamma}\left[\hat{t}(\gamma)\right]=\mathbb{E}_{\gamma,z}\left[\hat{q}(\gamma,z)\cdot\left(v(\gamma,z)-v_{\gamma}(\gamma,z)\frac{1-G(\gamma)}{g(\gamma)}\right)\right].$
@@
$\mathbb{E}_{\gamma}\left[\hat{t}(\gamma)\right]=\mathbb{E}_{\gamma,\theta}\left[q(\gamma,\theta)\cdot\left(\theta-\mathbf{V}(\theta|\gamma)\frac{1-G(\gamma)}{g(\gamma)}\right)\right].$
@@
PROOF OF PROPOSITION 1
Proof.
The proof proceeds in two steps. First, we show that under independence, it is without loss to consider allocation rules that are monotone in each dimension. Second, higher values of $\gamma$ correspond to pointwise higher values of $q$ when solving the problem $\gamma$-by-$\gamma$.
Lemma 3.
For every cyclically monotone $q(\gamma,\theta)$, there exists $\{\hat{q}^{j}(\gamma,\theta^{j})\}$ such that $q$ and $\{\hat{q}^{j}(\gamma,\theta^{j})\}$ produce the same surplus. Conversely, every collection $\{\hat{q}^{j}\}$ produces a cyclically monotone $q$ when concatenated.
Proof.
Suppose $q$ is cyclically monotone. We have that firm profits are
$\mathbb{E}_{\gamma,\theta}\left[\sum_{j}q^{j}(\gamma,\theta)\phi^{j}(\gamma,\theta^{j})\right]$ $=\sum_{j}\mathbb{E}_{\gamma,\theta}\left[q^{j}(\gamma,\theta)\phi^{j}(\gamma,\theta^{j})\right]$
$=\sum_{j}\mathbb{E}_{\gamma,\theta^{j}}\left[\mathbb{E}_{\theta^{-j}|\gamma}\left[q^{j}(\gamma,\theta)\right]\phi^{j}(\gamma,\theta^{j})\right]$
so let
$\hat{q}^{j}(\gamma,\theta^{j})=\mathbb{E}_{\theta^{-j}|\gamma}\left[q^{j}(\gamma,\theta)\right]=\int_{\Theta^{-j}}q^{j}(\gamma,\theta)\prod_{k\neq j}f^{k}(\theta^{k}|\gamma)d\theta^{-j}.$
Let $\hat{\theta}^{j}>\theta^{j}$; we have that
$\hat{q}^{j}(\gamma,\hat{\theta}^{j})-\hat{q}^{j}(\gamma,\theta^{j})=\int_{\Theta^{-j}}\left(q^{j}(\gamma,\hat{\theta}^{j},\theta^{-j})-q^{j}(\gamma,\theta^{j},\theta^{-j})\right)\prod_{k\neq j}f^{k}(\theta^{k}|\gamma)d\theta^{-j}.$
Applying cyclic monotonicity of $q$ to the two-cycle $(\hat{\theta}^{j},\theta^{-j});(\theta^{j},\theta^{-j})$ for arbitrary $\theta^{-j}$ gives that
$q^{j}(\gamma,\hat{\theta}^{j},\theta^{-j})\geq q^{j}(\gamma,\theta^{j},\theta^{-j})\implies q^{j}(\gamma,\hat{\theta}^{j},\theta^{-j})-q^{j}(\gamma,\theta^{j},\theta^{-j})\geq 0$
for all $\theta^{-j}$. Thus, integrating over $\Theta^{-j}$ gives that $\hat{q}^{j}(\gamma,\theta^{j})$ is monotone in $\theta^{j}$. By construction, $\{\hat{q}^{j}(\gamma,\theta^{j})\}$ produces the same surplus as $q$. The second statement of the Lemma is standard. ∎
By the Lemma, it suffice to optimize over $\{q^{j}(\gamma,\theta^{j})\}$ subject to (standard) monotonicity in $\theta^{j}$ opposed to cyclically monotone $q$. Let
$W^{j}(\gamma,\theta^{j})=\theta^{j}f^{j}(\theta^{j}|\gamma)g(\gamma)+F_{\gamma}(\theta^{j}|\gamma)G(\gamma)$
so the regularity in $\gamma$ assumption requires $W^{j}$ to be increasing in $\gamma$. The firm’s objective can be written as
$\max_{\{q^{j}\}}$ $\sum_{j}\int_{\Gamma}\int_{\Theta^{j}}q^{j}(\gamma,\theta^{j})W^{j}(\gamma,\theta^{j})d\theta^{j}d\gamma$
s.t. $\>q^{j}(\gamma,\theta^{j})\text{ increasing in }\theta^{j};$
$\text{ IC1}.$
Note here that analytically solving the problem in each dimension would require ironing. However, the proof of the proposition will not do any ironing; we will directly characterize the solution using monotone comparative statics.
Fix $j$. Monotone functions $q^{j}(\gamma,\cdot):\Theta^{j}\to[0,1]$ form a lattice under pointwise maximum/minimum. The set $\Gamma$ is straightforwardly a lattice. It suffice to show that
$O(q,\gamma)=\int_{\Theta^{j}}q(\theta^{j})W^{j}(\gamma,\theta^{j})d\theta^{j}$
is supermodular in $q,\gamma$. Fix $q,q^{\prime};\gamma,\gamma^{\prime}$. Without loss of generality, suppose $\gamma\geq\gamma^{\prime}$ since $\Gamma$ is one-dimensional. Note that for any $a,b,\max\{a,b\}=a+b-\min\{a,b\}$. We have
$O(\max\{q,q^{\prime}\},\max\{\gamma,\gamma^{\prime}\})+O(\min\{q,q^{\prime}\},\min\{\gamma,\gamma^{\prime}\})$
$=$ $\int_{\Theta^{j}}\max\{q(\theta^{j}),q^{\prime}(\theta^{j})\}W^{j}(\gamma,\theta^{j})d\theta^{j}+\int_{\Theta^{j}}\min\{q(\theta^{j}),q^{\prime}(\theta^{j})\}W^{j}(\gamma^{\prime},\theta^{j})d\theta^{j}$
$=$ $\int_{\Theta^{j}}\left[q(\theta^{j})+q^{\prime}(\theta^{j})-\min\{q(\theta^{j}),q^{\prime}(\theta^{j})\}\right]W^{j}(\gamma,\theta^{j})d\theta^{j}+\int_{\Theta^{j}}\min\{q(\theta^{j}),q^{\prime}(\theta^{j})\}W^{j}(\gamma^{\prime},\theta^{j})d\theta^{j}$
$=$ $\int_{\Theta^{j}}q(\theta^{j})W(\gamma,\theta^{j})d\theta^{j}+\int_{\Theta^{j}}\left[q^{\prime}(\theta^{j})-\min\{q(\theta^{j}),q^{\prime}(\theta^{j})\}\right]W^{j}(\gamma,\theta^{j})d\theta^{j}$
$+\int_{\Theta^{j}}\min\{q(\theta^{j}),q^{\prime}(\theta^{j})\}W^{j}(\gamma^{\prime},\theta^{j})d\theta^{j}$
$\geq$ $\int_{\Theta^{j}}q(\theta^{j})W(\gamma,\theta^{j})d\theta^{j}+\int_{\Theta^{j}}\left[q^{\prime}(\theta^{j})-\min\{q(\theta^{j}),q^{\prime}(\theta^{j})\}\right]W^{j}(\gamma^{\prime},\theta^{j})d\theta^{j}$
$+\int_{\Theta^{j}}\min\{q(\theta^{j}),q^{\prime}(\theta^{j})\}W^{j}(\gamma^{\prime},\theta^{j})d\theta^{j}$
$=$ $\int_{\Theta^{j}}q(\theta^{j})W(\gamma,\theta^{j})d\theta^{j}+\int_{\Theta^{j}}q^{\prime}(\theta^{j})W^{j}(\gamma^{\prime},\theta^{j})d\theta^{j}$
$=$ $O(q,\gamma)+O(q^{\prime},\gamma^{\prime})$
as desired. The inequity comes from the fact that pointwise $\left[q^{\prime}(\theta^{j})-\min\{q(\theta^{j}),q^{\prime}(\theta^{j})\}\right]\geq 0$ so $W$ being increasing in $\gamma$ gives that replacing $W^{j}(\gamma,\theta^{j})$ with $W^{j}(\gamma^{\prime},\theta^{j})$ makes the value smaller. ∎